# Taming OOD Actions for Offline Reinforcement Learning: An Advantage-Based Approach

### v3

@@ 5.2 Visualizing OOD Action Selection in PointMaze
In the previous subsection, we demonstrated that our method achieves state-of-the-art performance across a wide range of D4RL benchmark tasks, with particularly large gains on challenging sparse-reward environments. This improvement can be largely attributed to our newly designed advantage function, which enables the selection of beneficial OOD actions—a capability especially critical in sparse-reward tasks, where exploiting such actions is essential for high returns.
@@
To provide further visual evidence that the advantage function guides the selection of beneficial OOD actions, we conduct a comparative experiment on PointMaze. Specifically, we construct a toy environment based on the latest gymnasium-robotics implementation of PointMaze, derived from the Maze2D environment in the D4RL suite. As shown in Figure 4, the green circle indicates the agent’s starting position, the red star denotes the goal, and the beige squares represent static obstacles. The task involves navigating a 2-DoF point agent through a maze with obstacles to a fixed goal using Cartesian (x, y) actuation. This is a sparse-reward task: the agent receives a reward of 1 only upon reaching the goal and zero at all other steps.
We manually collected 853 successful yet sub-optimal trajectories to construct the offline training dataset, each shorter than 1000 steps but significantly longer than the shortest length of approximately 142 steps, which corresponds to a straight-line path from the start to the goal. As shown in the last subfigure of Figure 4, the offline dataset includes no optimal trajectories (length $<$ 175), only 3.4% near-optimal (length $<$ 225), and 19.1% competitive trajectories (length $<$ 300). These trajectories fall into three distinct modes—Left, Middle, and Right routes—depending on which corridor the agent takes to bypass the obstacles (see corresponding subfigures in Figure 4). The dataset comprises 391 391 Q-learning-style tuples of the form $(\bm{s},\bm{a},\bm{s^{\prime}},r,\text{done})$. The collection strategy and detailed trajectory quality analysis are provided in Appendix C.8.
@@
We train both DQL wang2022diffusion and ADAC for 50 000 steps using the constructed dataset to obtain their respective policies. Each policy is then evaluated in our environment, generating 300 trajectories per method, as illustrated in the bottom row of Figure 4.
@@
Compared to the full set of offline trajectories, the ADAC-generated trajectories contain a substantial number of straight-line paths from the start to the goal—routes that are entirely absent in the dataset. This visual evidence strongly supports that ADAC is capable of identifying and selecting superior OOD actions. In contrast, the trajectories generated by DQL, while showing moderate improvement over the offline data, still heavily reflect left, middle, and right patterns, indicating strong behavior cloning tendencies. Since the key distinction between DQL and ADAC lies in the introduction of advantage modulation, these visualizations clearly validate the effectiveness of the advantage function in enabling the selection of beneficial OOD actions. The last subfigure of Figure 4 provides quantitative evidence that ADAC substantially outperforms DQL in trajectory quality.
5.3 Ablation Study
One potential limitation of our method is the need to sample multiple actions ($N\equiv 25$ in this paper) from the behavior policy to compute the advantage function. However, our ablation study shows that in each task domain, competitive performance can be achieved across at least three different values of $\kappa$, indicating that the algorithm is relatively insensitive to $\kappa$. This suggests that the number of candidate actions can be reduced without significantly affecting performance. We implement our algorithm using the JAX/Flax framework, which offers faster training and inference speed than the DQL method and is comparable to the optimized EDP kang2023efficient implementation as shown in Figure 9.
@@ Appendix A Related Work
Offline RL focuses on learning effective policies solely from a pre-collected behavior dataset and has demonstrated significant success in practical applications [29, 31, 24]. The existing literature on offline RL can be classified into four main categories:
@@
Pessimistic value-based methods achieve conservatism by incorporating penalty terms into the value optimization objective, discouraging the value function from being overly optimistic on out-of-distribution (OOD) actions. Specifically, CQL [21] applies equal penalization to Q-values for all OOD samples, whereas EDAC [3] and PBRL [4] adjust the penalization based on the uncertainty level of the Q-value, measured using a neural network ensemble.
@@
Regularized policy-based methods constrain the learned policy to stay close to the behavior policy, thereby avoiding OOD actions. For instance, BEAR [20] constrains the optimized policy by minimizing the MMD distance to the behavior policy. BCQ [11] restricts the action space to those present in the dataset by utilizing a learned Conditional-VAE (CVAE) behavior-cloning model. Alternatively, TD3+BC [10] simply adds a behavioral cloning regularization term to the policy optimization objective and achieves excellent performance across various tasks. IQL [19] adopts an advantage-weighted behavior cloning approach, learning Q-value functions directly from the dataset. Meanwhile, DQL [36] leverages diffusion policies as an expressive policy class to enhance behavior-cloning. Our work falls into this category as we also incorporate a behavior-cloning term.
@@
Conditional sequence modeling methods induce conservatism by limiting the policy to replicate behaviors from the offline dataset [7, 38]. This approach naturally follows a supervised learning paradigm. Additionally, trajectories can be formulated as conditioned generative models and generated by diffusion models that satisfy conditioned constraints [14, 2].
@@
Model-based methods incorporate conservatism to prevent the policy from overgeneralizing to regions where the dynamics model predictions are unreliable. For example, COMBO [39] extends CQL to a model-based setting by enforcing small Q-values for OOD samples generated by the dynamics model. RAMBO [30] incorporates conservatism by adversarially training the dynamics model to minimize the value function while maintaining accurate transition predictions. Most model-based methods achieve conservatism through uncertainty quantification, penalizing rewards in regions with high uncertainty. Specifically, MOPO [40] uses the max-aleatoric uncertainty quantifier, MOReL [18] employs the max-pairwise-diff uncertainty quantifier, and MOBILE [34] leverages the Model-Bellman inconsistency uncertainty quantifier. Recently, [8] achieves conservatism by incorporating the value function inconsistency loss, enabling the training of a more reliable model.
@@ Appendix B Proof of Propositions
In this subsection, we provide comprehensive and complete proofs for our propositions listed in Section 3.
@@
$\frac{\partial\mathcal{L}(V)}{\partial V(\bm{s})}=$ $\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\bigg{[}\frac{\partial L_{2}^{\tau}(\delta)}{\partial V(\bm{s})}\bigg{]}$
=\ = $\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\bigg{[}\frac{\partial L_{2}^{\tau}(\delta)}{\partial\delta}\cdot\frac{\partial\delta}{\partial V(\bm{s})}\bigg{]}$
=\ = $\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\big{[}2|1-\mathds{1}(\delta<0)|\delta(\bm{s},\bm{a},\bm{s^{\prime}})\cdot(-1)\big{]},$
@@
From Lemma 1 in [19], which states that
@@
We first show that $\|A(\bm{a^{\prime}}|\bm{s^{\prime}})\|_{\infty}\leq 2R_{\rm max}/(1-\gamma)$. From the proof of Proposition 3.2, we know that
@@ C.1 Advantage Function Characterization and Regularization
halfcheetah-medium $1.62_{\pm 0.2}$ $-0.11_{\pm 0.3}$ 32.8
halfcheetah-medium-replay $1.84_{\pm 0.3}$ $-0.90_{\pm 0.0}$ 30.4
halfcheetah-medium-expert $2.00_{\pm 0.2}$ $-1.75_{\pm 0.3}$ 38.7
hopper-medium $0.48_{\pm 0.1}$ $-0.14_{\pm 0.0}$ 22.1
hopper-medium-replay $1.25_{\pm 0.2}$ $-0.83_{\pm 0.1}$ 38.3
hopper-medium-expert $0.39_{\pm 0.2}$ $-0.55_{\pm 0.0}$ 10.1
walker2d-medium $0.70_{\pm 0.1}$ $-0.06_{\pm 0.0}$ 43.5
walker2d-medium-replay $1.87_{\pm 0.3}$ $-0.13_{\pm 0.0}$ 20.0
walker2d-medium-expert $2.33_{\pm 0.3}$ $-2.20_{\pm 0.5}$ 26.2
antmaze-umaze $1.70_{\pm 0.2}$ $-1.05_{\pm 0.1}$ 52.8
antmaze-umaze-diverse $0.03_{\pm 0.0}$ $-0.04_{\pm 0.0}$ 34.0
antmaze-medium-play $0.58_{\pm 0.1}$ $-0.40_{\pm 0.1}$ 44.7
antmaze-medium-diverse $0.48_{\pm 0.05}$ $-0.26_{\pm 0.05}$ 54.0
antmaze-large-play $0.58_{\pm 0.02}$ $-0.28_{\pm 0.05}$ 48.3
antmaze-large-diverse $0.42_{\pm 0.08}$ $-0.23_{\pm 0.02}$ 62.3
pen-human $2.34_{\pm 1.1}$ $-1.74_{\pm 0.6}$ 41.3
pen-cloned $1.14_{\pm 0.2}$ $-1.01_{\pm 0.1}$ 33.4
kitchen-complete $1.15_{\pm 0.2}$ $-0.85_{\pm 0.1}$ 42.3
kitchen-partial $0.47_{\pm 0.1}$ $-0.44_{\pm 0.2}$ 31.1
kitchen-mixed $0.52_{\pm 0.0}$ $-0.74_{\pm 0.0}$ 33.8
@@ C.3 Repeated Sampling for Value-Guided Diffusion Policies
During critic updates, we widely employ a Max-Q Backup mechanism from CQL [21]: for each transition, multiple candidate actions are sampled from the policy at the next state, and the Q-target is computed using the maximum or softmax-weighted Q-value among them. This mitigates underestimation bias caused by poor single-sample rollouts and reduces the variance of the TD targets. We observe that modest sample counts (e.g., 3–5) already improve stability, while more complex tasks—such as halfcheetah-medium-replay and antmaze-medium-diverse benefit from larger sample sizes (e.g., 10). As shown in Figure 8, increasing the number of backup samples leads to higher predicted Q-values and improved empirical returns.
@@
At evaluation time, we sample a large set of candidate actions (typically 50–200) from the diffusion model, and select actions via Q-weighted sampling. Specifically, Q-values are transformed into a softmax distribution, from which the final action is drawn. This Q-guided inference biases the policy toward high-value modes while retaining stochasticity, offering a principled alternative to greedy or deterministic selection. Across tasks, we consistently observe superior returns compared to single-sample decoding.
@@ C.4 Hyperparameter Setup
We highlight several key components of our approach, including the use of the $\kappa$ quantile to define reward thresholds, the integration of residual blocks to enhance the expressiveness of the critic, and repeated sampling to exploit the multimodal capacity of the diffusion model. These components play a central role in achieving strong performance across tasks. Table 3 provides the full set of hyperparameters; all other parameters follow default configurations from DQL [36] without further modification.
@@
Task $\boldsymbol{\kappa}$ Max Q Backup $\boldsymbol{n}$ Critic Net
@@
C.5 Alternative Implementation
While expectile regression provides a theoretically grounded approach for estimating the optimal value function from offline data, we identify an IQL-style alternative that achieves comparable empirical performance and is easier to tune, particularly under highly suboptimal dataset quality and coverage. This approach couples value estimation with policy optimization, leveraging the evolving policy to gradually shift the data distribution toward regions with more informative value estimates. As a result, it often makes hyperparameter tuning, such as selecting $\tau$, more robust in practice.
Algorithm 2 Alternative Impl. for ADAC
1:  Input initial all parameters
2:  Train a diffusion behavior policy
3:  Train a MLP transition model $P$
4:  for each iteration do
5:     Update $\mathcal{L}_{v}(\varphi)$ according to Eq. (16)
6:     Obtain $A(\bm{a^{\prime}}|\bm{s^{\prime}})$ according to Eq. (9)
7:     Update $\mathcal{L}_{q}(\phi)$ according to Eq. (14)
8:     Update $\mathcal{L}_{p}(\theta)$ according to Eq. (15)
9:     Soft update parameter $\phi$ and $\theta$
10:  end for
Building on this observation, we replace direct value regression with a IQL-style target that fits $V$ to the $\tau$-expectile of the critic $Q_{\phi}$:
$\mathcal{L}_{v}(\varphi)=\mathbb{E}_{\bm{s}\sim\mathcal{D}}\!\Bigl{[}L_{2}^{\tau}\!\bigl{(}Q_{\phi}(\bm{s},\bm{a})-V_{\varphi}(\bm{s})\bigr{)}\Bigr{]}.$ (16)
This formulation promotes stable learning by anchoring $V$ to $Q_{\phi}$ and leveraging the advantage-modulated Q-function to improve value estimation. The full procedure is summarized in Algorithm 2. Empirically, this variant achieves value estimates comparable to the original design and demonstrates substantial robustness to the choice of $\tau$ on typical suboptimal offline datasets.
C.6 Computational Efficiency of Training and Inference
Figure 9: Speedup comparison of ADAC, EDP and DQL baseline methods.
All experiments were conducted using NVIDIA RTX 4090 GPUs. Our implementation builds on jaxrl_m, a modular and extensible offline RL codebase based on Flax and JAX. We extend this framework to the proposed Advantage-Based Diffusion Actor-Critic (ADAC).
The optimized ADAC method achieves a training throughput of approximately 115 steps per second, resulting in a 2×2$\times$2 × speedup during training and a 1.64×1.64$\times$1.64 × speedup in inference compared to the original DQL implementation, as shown in Figure 9. ADAC matches the training efficiency of Efficient Diffusion Policy (EDP) [17] while offering superior inference speed. These results highlight the computational efficiency of our approach in both training and inference.
C.7 Auxiliary Model Pretraining

### v4

@@ 5.2 Visualizing OOD Action Selection in PointMaze
In the previous subsection, we demonstrated that our method achieves SOTA performance across a wide range of D4RL benchmark tasks, with particularly large gains on challenging sparse-reward environments. This improvement can be largely attributed to our newly designed advantage function, which enables the selection of beneficial OOD actions—a capability especially critical in sparse-reward tasks.
@@
To better visualize the strength of ADAC in guiding the selection of beneficial OOD actions, we conduct a comparative experiment on PointMaze. Specifically, we construct a toy environment based on the latest gymnasium-robotics (de Lazcano et al., 2023) implementation of PointMaze, derived from the Maze2D environment in the D4RL suite. As shown in Figure˜4, the green circle indicates the agent’s starting position, the red star denotes the goal, and the beige squares represent static obstacles. The task involves navigating a 2-DoF point agent through a maze with obstacles to a fixed goal using Cartesian $(x,y)$ actuation. This is a sparse-reward task: the agent receives a reward of 1 only upon reaching the goal and zero at all other steps. We manually collect 853 trajectories of varying quality, as illustrated in the bar plot of Figure 4, which together yield 391,391 tuples of the form $(\bm{s},\bm{a},\bm{s^{\prime}},r,\text{done})$. Details of the dataset collection strategy and the trajectory quality analysis are provided in Appendix C.7.
@@
We train both DQL (Wang et al., 2022) and ADAC for 50 000 steps using the constructed dataset to obtain their respective policies. Each policy is then evaluated in our environment, generating 300 trajectories per method, as illustrated in the bottom row of Figure 4.
@@
Note that the original collected offline trajectories does not contain any optimal trajectories (of steps less than 175, see the bar plot of Figure 4). Nevertheless, ADAC effectively learns the optimal trajectories from these suboptimal datasets, which generated a substantial number of straight-line (optimal) trajectories from the start to the goal—routes that are entirely absent in the dataset. This visual evidence strongly supports that ADAC is capable of identifying and selecting superior OOD actions. In contrast, the trajectories generated by DQL, while showing moderate improvement over the offline data, still largely follow left-, middle-, and right-pattern behaviors, indicating a strong tendency toward behavior cloning. Since the key distinction between DQL and ADAC lies in the introduction of advantage modulation, these visualizations clearly validate the effectiveness of the advantage function in enabling the selection of beneficial OOD actions. The last subfigure of Figure˜4 provides quantitative evidence that ADAC substantially outperforms DQL in trajectory quality.
Figure 5: Ablation of the advantage component across four domains. Bars show domain-level mean normalized scores.
5.3 Ablation on the Advantage Component
We assess the importance of the advantage by comparing models with and without this component on four representative domains: Gym Locomotion, AntMaze, Adroit, and Kitchen. The ablation results in Figure˜5 show that introducing the advantage consistently strengthens performance across all settings, with relative improvements of 11.1% on Gym Locomotion, 12.2% on AntMaze, 10.9% on Adroit, and 12.4% on Kitchen. Beyond the overall gains, the pattern is consistent across domains with diverse dynamics and reward structures, indicating that the advantage contributes broadly rather than acting as a domain-specific tweak.
Figure 6: Training and inference speedups of ADAC compared to EDP and the DQL baseline.
5.4 Computational Efficiency of Training and Inference
All experiments were conducted on a single NVIDIA RTX 4090, without any distributed training or model parallelism. This setup ensures that reported throughput reflects algorithmic and implementation efficiency rather than scale-out effects.
Building on the lightweight utilities of jaxrl_m (Flax/JAX), we reimplemented Diffusion QL in JAX and subsequently introduced advantage-centric components on top of it. The resulting codebase follows a flat, modular design that facilitates reproduction and portability across tasks. Under identical hardware and evaluation protocols, the implementation delivers consistent throughput, yielding a $2\times$ improvement in training and a $1.64\times$ improvement in inference over the original DQL implementation. In head-to-head comparisons with Efficient Diffusion Policy (EDP) (Kang et al., 2023), training throughput is comparable, while inference is faster (Fig. 6).
@@ Appendix A Related Work
Offline RL focuses on learning effective policies solely from a pre-collected behavior dataset and has demonstrated significant success in practical applications (Rafailov et al., 2021; Singh et al., 2020; Li et al., 2010). The existing literature on offline RL can be classified into four main categories:
@@
Pessimistic value-based methods achieve conservatism by incorporating penalty terms into the value optimization objective, discouraging the value function from being overly optimistic on out-of-distribution (OOD) actions. Specifically, CQL (Kumar et al., 2020) applies equal penalization to Q-values for all OOD samples, whereas EDAC (An et al., 2021) and PBRL (Bai et al., 2022) adjust the penalization based on the uncertainty level of the Q-value, measured using a neural network ensemble.
@@
Regularized policy-based methods constrain the learned policy to stay close to the behavior policy, thereby avoiding OOD actions. For instance, BEAR (Kumar et al., 2019) constrains the optimized policy by minimizing the MMD distance to the behavior policy. BCQ (Fujimoto et al., 2019) restricts the action space to those present in the dataset by utilizing a learned Conditional-VAE (CVAE) behavior-cloning model. Alternatively, TD3+BC (Fujimoto & Gu, 2021) simply adds a behavioral cloning regularization term to the policy optimization objective and achieves excellent performance across various tasks. IQL (Kostrikov et al., 2021) adopts an advantage-weighted behavior cloning approach, learning Q-value functions directly from the dataset. Meanwhile, DQL (Wang et al., 2022) leverages diffusion policies as an expressive policy class to enhance behavior-cloning. Our work falls into this category as we also incorporate a behavior-cloning term.
@@
Conditional sequence modeling methods induce conservatism by limiting the policy to replicate behaviors from the offline dataset (Chen et al., 2021; Wu et al., 2023). This leads to a supervised learning paradigm. Additionally, trajectories can be formulated as conditioned generative models and generated by diffusion models that satisfy conditioned constraints (Janner et al., 2022; Ajay et al., 2022).
@@
Model-based methods incorporate conservatism to prevent the policy from overgeneralizing to regions where the dynamics model predictions are unreliable. For example, COMBO (Yu et al., 2021) extends CQL to a model-based setting by enforcing small Q-values for OOD samples generated by the dynamics model. RAMBO (Rigter et al., 2022) incorporates conservatism by adversarially training the dynamics model to minimize the value function while maintaining accurate transition predictions. Most model-based methods achieve conservatism through uncertainty quantification, penalizing rewards in regions with high uncertainty. Specifically, MOPO (Yu et al., 2020) uses the max-aleatoric uncertainty quantifier, MOReL (Kidambi et al., 2020) employs the max-pairwise-diff uncertainty quantifier, and MOBILE (Sun et al., 2023) leverages the Model-Bellman inconsistency uncertainty quantifier. Recently, (Chen et al., 2025) achieves conservatism by incorporating the value function inconsistency loss, enabling the training of a more reliable model.
@@ Appendix B Proof of Propositions
In this subsection, we provide comprehensive and complete proofs for our propositions listed in Section˜3.
@@
$\frac{\partial\mathcal{L}(V)}{\partial V(\bm{s})}=\$$\mathbb{E}_{\bm{a}\sim\mu(\cdot|\bm{s}),\bm{s^{\prime}}\sim P(\cdot|\bm{s},\bm{a})}\bigg[\frac{\partial L_{2}^{\tau}(\delta)}{\partial V(\bm{s})}\bigg]$
[…]
@@
From Lemma 1 in (Kostrikov et al., 2021), which states that
@@
We first show that $\|A(\bm{a^{\prime}}|\bm{s^{\prime}})\|_{\infty}\leq 2R_{\rm max}/(1-\gamma)$. From the proof of Proposition˜3.2, we know that
@@ C.1 Advantage Function Characterization and Regularization
halfcheetah-medium $1.62_{\pm 0.2}$ $0.11_{\pm 0.3}$ 32.8
halfcheetah-medium-replay $1.84_{\pm 0.3}$ $0.90_{\pm 0.0}$ 30.4
halfcheetah-medium-expert $2.00_{\pm 0.2}$ $1.75_{\pm 0.3}$ 38.7
hopper-medium $0.48_{\pm 0.1}$ $0.14_{\pm 0.0}$ 22.1
hopper-medium-replay $1.25_{\pm 0.2}$ $0.83_{\pm 0.1}$ 38.3
hopper-medium-expert $0.39_{\pm 0.2}$ $0.55_{\pm 0.0}$ 10.1
walker2d-medium $70_{\pm 0.1} 0.0$$06_{\pm 0.0}$ 43.5
walker2d-medium-replay $1.87_{\pm 0.3}$ $0.13_{\pm 0.0}$ 20.0
walker2d-medium-expert $33_{\pm 0.3} 2.2$$20_{\pm 0.5}$ 26.2
antmaze-umaze $1.70_{\pm 0.2}$ $1.05_{\pm 0.1}$ 52.8
antmaze-umaze-diverse $03_{\pm 0.0} 0.0$$04_{\pm 0.0}$ 34.0
antmaze-medium-play $0.58_{\pm 0.1}$ $0.40_{\pm 0.1}$ 44.7
antmaze-medium-diverse $0.48_{\pm 0.05}$ $0.26_{\pm 0.05}$ 54.0
antmaze-large-play $0.58_{\pm 0.02}$ $0.28_{\pm 0.05}$ 48.3
antmaze-large-diverse $0.42_{\pm 0.08}$ $0.23_{\pm 0.02}$ 62.3
pen-human $2.34_{\pm 1.1}$ $1.74_{\pm 0.6}$ 41.3
pen-cloned $1.14_{\pm 0.2}$ $1.01_{\pm 0.1}$ 33.4
kitchen-complete $1.15_{\pm 0.2}$ $0.85_{\pm 0.1}$ 42.3
kitchen-partial $7_{\pm 0.1} 0.44±0$$4_{\pm 0.2}$ 31.1
kitchen-mixed $0.52_{\pm 0.0}$ $0.74_{\pm 0.0}$ 33.8
@@ C.3 Repeated Sampling for Value-Guided Diffusion Policies
During critic updates, we widely employ a Max-Q Backup mechanism from CQL (Kumar et al., 2020): for each transition, multiple candidate actions are sampled from the policy at the next state, and the Q-target is computed using the maximum or softmax-weighted Q-value among them. This mitigates underestimation bias caused by poor single-sample rollouts and reduces the variance of the TD targets. We observe that modest sample counts (e.g., 3–5) already improve stability, while more complex tasks—such as halfcheetah-medium-replay and antmaze-medium-diverse benefit from larger sample sizes (e.g., 10). As shown in Figure 9, increasing the number of backup samples leads to higher predicted Q-values and improved empirical returns.
@@
At evaluation time, we sample a large set of candidate actions (typically 50–200) from the diffusion model, and select actions via Q-weighted sampling. Specifically, Q-values are transformed into a softmax distribution, from which the final action is drawn. This Q-guided inference biases the policy toward high-value modes while retaining stochasticity. Across tasks, we consistently observe superior returns compared to single-sample decoding.
@@ C.4 Hyperparameter Setup
We highlight several key components of our approach, including the use of the $\kappa$ quantile to define reward thresholds, the integration of residual blocks to enhance the expressiveness of the critic, and repeated sampling to exploit the multimodal capacity of the diffusion model. These components play a central role in achieving strong performance across tasks. Table 3 provides the full set of hyperparameters; all other parameters follow default configurations from DQL (Wang et al., 2022) without further modification.
@@
Task $\bm{\kappa}$ Max Q Backup $\bm{n}$ Critic Net
@@
C.5 Sensitivity Analysis of the Parameter $\kappa$
One potential limitation of our method is the need to sample multiple actions ($N\equiv 25$ in this paper) from the behavior policy to compute the advantage function. However, our ablation study shows that in each task domain, competitive performance can be achieved across at least three different values of $\kappa$, indicating that the algorithm is relatively insensitive to $\kappa$. This suggests that the number of candidate actions can be reduced without significantly affecting performance. We implement our algorithm using the JAX/Flax framework, which offers faster training and inference speed than the DQL method and is comparable to the optimized EDP (Kang et al., 2023) implementation as shown in Figure˜6.
C.6 Auxiliary Model Pretraining